OpenStudy (anonymous):

What is the S to the 10-th power, if S is the sum of the solutions of the equation 2x2 - 2ix +10=0, where i=√(-1).

hartnn (hartnn):

sum of roots = -b/a in ax^2+bx+c=0

hartnn (hartnn):

here sum = 2i/2 = i so i^10=?

OpenStudy (anonymous):

-1

hartnn (hartnn):

yup